# Fixed-domain curve counts for blow-ups of projective space

### v3

@@ 2. Step 2
    For any $p,j$ as above, let $\alpha_{p,j}>0$ be the maximal order to which both properties (i), (ii) hold, that is, $D_{\{j\}}$ contains $p$ with multiplicity $\alpha_{p,j}$ and $f_{j}$ vanishes at $p$ to order $\alpha_{p,j}$. Then, make the following modifications (independently for all $j$):
@@
    • $D_{\{j\}}$ is replaced by $D_{\{j\}}-\alpha_{p,j}p$, while all $D_{S}$ with #$\#S>1$ remain empty, and
@@
    After step 2, we have $\mathcal{L},D_{S},f_{j}$ (still $D_{S}=0$ if #$\#S>1$) with the property that, for any $p\in C^{\circ}$ and any $j\in[r+1]$, at most one of the properties (i), (ii) hold.
@@ 3. Step 3
    As a section of $\mathcal{L}$, the order of vanishing of $f_{j}$ at $p$ is at least $\alpha_{\text{tot}}-\alpha_{j}$, and, after step 2, we have equality whenever $j\in S_{1}$. In particular, the common order of vanishing of the $f_{j}$ is equal to $\alpha_{\text{tot}}-\alpha_{\max}$, which is strictly positive exactly when #$\#S_{1}\geq 2$.
@@
    $[r+1]=S^{1}\supsetneq S^{2}\supsetneq\cdots\supsetneq S^{t}$
@@
    • each $D_{\{j\}}$ is (temporarily) replaced by $D_{\{j\}}-\alpha_{j}p$,
@@ 4. Step 4
    This step is identical to step 1, with $p_{i}$ in place of $p$, and the $D_{S}$ with #$\#S>1$ playing no role. After step 4, it will no longer be true that $p_{i}$ satisfies (ii) for all $j$.
@@ 5. Step 5
    This step is new. For a fixed $p_{i}$, $\alpha>0$ be the largest order to which $p_{i}$ is contained in all of the $D_{j}$ simultaneously. Then, make the following modifications:
    • $\mathcal{L}$ is replaced by $\mathcal{L}((-\alpha r)p)$,
    • each $D_{\{j\}}$ is replaced by $D_{\{j\}}-\alpha p$, while all $D_{S}$ with #$\#S>1$ remain unchanged, and
@@
    The final two steps are only needed for wild and inactive $p_{i}$.
@@ 7. Step 7
    Repeat step 3 with $p_{i}$ in place of $p$. Here, the subset $S^{i}_{\circ,1}$ plays the role of $S_{1}$ in step 3 above. Note that, by Lemma 43, step $7$ is run only if step $6$ is.
@@
First, suppose that $p_{i}$ is inactive. As we have noted, the condition $V(x_{i})$ no longer imposes constraints on the $f_{j},D_{S}$ post-twisting. On the other hand, after steps 6 and 7, the divisors $D_{S}$ containing $p_{i}$ impose conditions on the intersection of $\text{im}(f)$ with the torus-invariant boundary of $\mathbb{P}^{r}$ at $p_{i}$ in exactly the same way as at the $p\in C^{\circ}$. Then, we summarize the situation:
@@
where delimited-⟨$\langle$-$\rangle$ - ⟩ denotes linear span (so that $\Lambda_{i}$ has dimension #$\#S^{i}_{\circ}$).
@@
In the unique lift $f_{s}:C\to Y_{s}$ of $f$ to $Y_{s}$, for each $S$ with #$\#S=s+1$, the divisor $D_{S}$ is constrained to map to strict transform of $\Lambda_{S}$ in $Y_{s}$. Therefore, the blow-up $\rho_{s}$ decreases the degree of $C$ by at least $(r-\#S)k_{S}=(r-s-1)k_{S}$ for each such $S$, that is,
@@
The claim now follows from the fact that, by construction, the sections $f_{j}\in H^{0}(C,\mathcal{L})$ comprising $f:C\to\mathbb{P}^{r}$ have the same order of vanishing for $j\in(S_{\circ}^{i})^{c}$. Therefore, twisting $\mathcal{L}$ down by this order of vanishing yields a point $y_{k^{\prime}}$ in the sequence $y=(y_{0},\ldots,y_{k})$ for which the corresponding coordinates are all non-zero and equal to each other, for the same reason as in Lemma 45. $\Diamond$
@@
The quantities $\deg(\mathcal{L}),\deg(D_{S})$ are taken to be those that are changing throughout the twisting algorithm. As for the last term, we somewhat abusively set the values of $(r-\#S^{i}_{\circ})$ to be equal to $r$ initially (as if all marked points $p_{i}$ are initially regular), so that the last summation is initially equal to $-rn$, and, by (1), $\Omega$ is initially equal to zero. Then, if $p_{i}$ is inactive, the summand $(r-\#S^{i}_{\circ})$ is removed after either step 4 or 5 (whichever is applied first). If $p_{i}$ is wild, then the value of #$\#S^{i}_{\circ}$ is set to the correct one after step 6. If $p_{i}$ is regular, the summand $(r-\#S^{i}_{\circ})$ remains equal to $r$ throughout. The quantities $\deg(\mathcal{L}),\deg(D_{S})$ are taken to be those that are changing throughout the twisting algorithm.
By Lemma 46, the final value of $\Omega$ is an upper bound for the expected relative dimension of $\tau^{\prime}$. Therefore, if $f_{\text{init}}\notin\mathbb{P}^{\circ}$, that is, at least one step of the algorithm is run, then the expected relative dimension of $\tau^{\prime}$ will be strictly negative, as needed.
@@ 3. Step 3
    First, decreasing $\deg(\mathcal{L})$ by $\alpha_{\text{tot}}-\alpha_{\max}$ decreases $\Omega$ by $(r+1)(\alpha_{\text{tot}}-\alpha_{\max})$, and replacing each $D_{\{j\}}$ by $D_{\{j\}}-\alpha_{j}p$ increases it by $(r-1)\alpha_{\text{tot}}$. Finally, the modification of $D_{[r+1]\backslash S^{m}}$ decreases $\Omega$ by (#$(\#S^{m}-1)(\alpha^{\prime}(m)-\alpha^{\prime}(m-1))$. We thus need to show that
    $-2\alpha_{\text{tot}}+(r+1)\alpha_{\max}-\sum_{m=2}^{t}(\#S^{m}-1)(\alpha^{\prime}(m)-\alpha^{\prime}(m-1))<0.$
@@ 5. Step 5
  • if $p_{i}$ is wild, then step 6 decreases $\deg(\mathcal{L})$ and $\deg(D_{\{j\}})$ together by least 1 for each $j\in S^{i}_{\circ}$, decreasing $\Omega$ by at least $2\cdot\#S^{i}_{\circ}$. Furthermore, the $-(r-\#S^{i}_{\circ})$ is updated to the correct value (from $-r$ originally) in this step. In total, the value of $\Omega$ goes down at least by
    $02\cdot\#S^{i}_{\circ}+(r-\#S^{i}_{\circ})-r>$ 0
    in step 6. In step 7, $\Omega$ can again only decrease, by step 3. (Note that step $7$ does not affect the contribution of $-(r-\#S^{i}_{\circ})$.)
@@
Therefore, the conclusion follows from Lemma 1.
$\Diamond$
Proof of Proposition 34.
Lemmas 49 and 50 at long last imply that $V(x_{1},\ldots,x_{n})$ is contained in $\mathbb{P}^{\circ}$. Indeed, if this is not the case for some $f\in V(x_{1},\ldots,x_{n})$, then applying the twisting algorithm to $f$ yields a point on ${\mathcal{M}}_{g,n}(Y,\bar{\beta})$ moving in a family dominating the target of $\tau^{\prime}$, but this is a contradiction due to dimension reasons.
@@
It remains to check that any $f\in V(x_{1},\ldots,x_{n})$ has no non-trivial tangent vectors. It is elementary to check that such a tangent vector would give rise to a non-trivial relative tangent vector of the map $\tau:{\mathcal{M}}_{g,n}(X,\beta)\to{\mathcal{M}}_{g,n}\times X^{n}$ over a general point, which contradicts Lemma 1. $\Diamond$
@@ 3.2. Integral formula on $S$
In this section, we push forward the formula of Theorem 10 to obtain Theorem 11. Up to setting $k_{i}=0$ for $i>\ell$, we can assume that $\ell=r+1$.
@@
~$\widetilde{\mathcal{E}}:=\mathcal{E}\otimes\mathcal{O}(N)=\bigoplus_{i=1}^{r+1}\nu_{*}(\mathcal{P}(-\mathcal{D}+N+\mathcal{D}_{i}))=\bigoplus_{i=1}^{r+1}\nu_{*}(\mathcal{P}(\Delta_{i})),$
@@
$N=\sum_{i=1}^{r+1}N_{i},\ \mathcal{D}=\sum_{i=1}^{r+1}\mathcal{D}_{i}\text{ and }\Delta_{i}=-\mathcal{D}+N+\mathcal{D}_{i}\text{ for }i=1,\ldots,r+1.$
@@
where in the last equality we have pushed forward to $S$. It remains to compute $c(\widetilde{\mathcal{E}})$; this will be an application of Grothendieck-Riemann-Roch.
@@
3.2.2. Computation of $c(\widetilde{\mathcal{E}})$
We now express $c(\widetilde{\mathcal{E}})$ in terms of the classes defined above. Fix an index $i$ with $11\leq i\leq r+$ 1.
Proposition 52.
@@ 3.2.2. Computation of $c(\widetilde{\mathcal{E}})$
where in the second equality we used identity (i) in Lemma 51. Similarly, we have
@@
Here, in the second equality we have used identity (ii) in Lemma 51. After summing all the obtained contributions, we obtain the claim. $\Diamond$
Corollary 53.
@@
Substituting 53 into (17) completes the proof of Theorem 11, namely that
@@ 3.3. Specialization to genus 0
We now specialize Theorem 11 to $g=0$ to prove Theorem 12.
@@
Then, the class $\eta_{i}$ is the first Chern class of $\mathcal{O}_{\mathbb{P}^{k_{i}}}(1)$ and Theorem 11 gives:
@@
This concludes the proof of Theorem 12.
@@ 3.4. Specialization to $X={\mathsf{Bl}}_{q}(\mathbb{P}^{r})$
Theorem 11 also admits a reasonably elegant specialization to $X={\mathsf{Bl}}_{q}(\mathbb{P}^{r})$. We take now $k_{2}=\ldots=k_{r+1}=0$ and write $k=k_{1}$, $\eta=\eta_{1}$, and $\mathsf{E}=\mathsf{E}_{1}$, so that $\tau=\overline{\tau}_{2}=\cdots=\overline{\tau}_{r+1}$ and $x=\overline{x}_{2}=\cdots=\overline{x}_{r+1}$.
@@
$=\mathrm{Coeff}\left(A(\eta)(C(\eta)(1+\eta B(\eta)-\eta D(\eta)^{2})^{g};\eta^{k}\right)$
@@
From Lemma 54, we obtain the residue formula
@@
(${\displaystyle{m+n\choose r}=\sum_{k=0}^{r}{m\choose k}{n\choose r-k}}\text{ for all }r,n,m\in\mathbb{Z}_{\geq 0}.$
Finally, replacing $a$ with $g-m$, we obtain Theorem 13.
@@ 4. Virtual counts for $X={\mathsf{Bl}}_{q}(\mathbb{P}^{r})$
In this section, we establish the virtual count of Theorem 14.
@@ 4.1. Preliminaries on $QH^{*}(X)$
Because $X$ is Fano, by [20, Proposition 2.2], we have
@@
as $\mathbb{Q}[\mathsf{C}]$-modules. Here, $\mathsf{C}\subseteq H_{2}(X,\mathbb{Z})$ is the cone of effective curves in $X$, and by [10, Proposition 4.1], we have
@@
are the only two primitive sets in the sense of [14, Definition 1.1] and, using [14, Theorem 1.2], from these two sets we respectively obtain the following two relations in $QH^{*}(X,\mathbb{Q})$:
@@
so the above collection spans $QH^{*}(X,\mathbb{Q})$. The fact that they are linearly independent follows from [14, Theorem 1.2]. $\Diamond$
@@
of $H^{*}(X,\mathbb{Q})$. We now express each of these classes in $QH^{*}(X,\mathbb{Q})$ as linear combinations of the basis elements of Lemma 55 using [14, Theorem 1.6]. In order for [14, Theorem 1.6] to apply, we need the following result:
Lemma 56.
@@
Immediate consequence of [14, Theorem 3.9]. $\Diamond$
Lemma 57.
@@
We start with $\mathsf{H}^{i}$. Let $\alpha_{i}$ be the cone generated by $v_{r},\ldots,v_{i+1}$ for $i=1,\ldots,r$. We have that $\alpha_{i}$ corresponds to a toric subvariety of class $\mathsf{H}^{r-i}$. Then, [14, Theorem 1.6] and the fact the only exceptional set special for $\alpha_{i}$ is the empty set yields the desired equality.

### v4

@@ 2. Step 2
    For any $p,j$ as above, let $\alpha_{p,j}>0$ be the maximal order to which both properties (i), (ii) hold, that is, $D_{\{j\}}$ contains $p$ with multiplicity $\alpha_{p,j}$ and $f_{j}$ vanishes at $p$ to order $\alpha_{p,j}$. Then, make the following modifications (independently for all $j$).
@@
    • $D_{\{j\}}$ is replaced by $D_{\{j\}}-\alpha_{p,j}p$, while all $D_{S}$ with #S>1\#S>1 remain empty, and
@@
    After step 2, we have $\mathcal{L},D_{S},f_{j}$ (still $D_{S}=0$ if #S>1\#S>1) with the property that, for any $p\in C^{\circ}$ and any $j\in[r+1]$, at most one of the properties (i), (ii) hold.
@@ 3. Step 3
    As a section of $\mathcal{L}$, the order of vanishing of $f_{j}$ at $p$ is at least $\alpha_{\text{tot}}-\alpha_{j}$, and, after step 2, we have equality whenever $j\in S_{1}$. In particular, the common order of vanishing of the $f_{j}$ is equal to $\alpha_{\text{tot}}-\alpha_{\max}$, which is strictly positive exactly when $S_{1}\geq 2$.
@@
    $r+1]=S^{1}\supsetneq S^{2}\supsetneq\cdots\supsetneq S^{t}$
@@
    • each $D_{\{j\}}$ is (temporarily, possibly to be modified further in the bullet point below) replaced by $D_{\{j\}}-\alpha_{j}p$,
@@ 4. Step 4
    This step is identical to step 1, with $p_{i}$ in place of $p$, and the $D_{S}$ with #S>1\#S>1 playing no role. After step 4, it will no longer be true that $p_{i}$ satisfies (ii) for all $j$.
@@ 5. Step 5
    This step is new. For a fixed $p_{i}$, $\alpha>0$ be the largest order to which $p_{i}$ is contained in all of the $D_{j}$ simultaneously. Then, make the following modifications.
    • $\mathcal{L}$ is replaced by $\mathcal{L}((-\alpha r)p_{i})$,
    • each $D_{\{j\}}$ is replaced by $D_{\{j\}}-\alpha p_{i}$, while all $D_{S}$ with #S>1\#S>1 remain unchanged, and
@@
    The final two steps are only needed for $p_{i}$ that are either inactive or wild.
@@ 7. Step 7
    Repeat step 3 with $p_{i}$ in place of $p$. Here, the subset $S^{i}_{\circ,1}$ plays the role of $S_{1}$ in step 3 above. Note that, by Lemma 3.11, step $7$ is run only if step $6$ is.
@@
First, suppose that $p_{i}$ is inactive. As we have noted, the condition $V(x_{i})$ no longer imposes constraints on the $f_{j},D_{S}$ post-twisting. On the other hand, after steps 6 and 7, the divisors $D_{S}$ containing $p_{i}$ impose conditions on the intersection of $\text{im}(f)$ with the torus-invariant boundary of $\mathbb{P}^{r}$ at $p_{i}$ in exactly the same way as at the $p\in C^{\circ}$. Then, we summarize the situation.
@@
where $\langle$-$\rangle$ denotes linear span (so that $\Lambda_{i}$ has dimension $S^{i}_{\circ}$).
@@
In the unique lift $f_{s}:C\to Y_{s}$ of $f$ to $Y_{s}$, for each $S$ with #S=s+1\#S=s+1, the divisor $D_{S}$ is constrained to map to strict transform of $\Lambda_{S}$ in $Y_{s}$. Therefore, the blow-up $\rho_{s}$ decreases the degree of $C$ by at least $(r-\#S)k_{S}=(r-s-1)k_{S}$ for each such $S$, that is,
@@
The claim now follows from the fact that, by construction, the sections $f_{j}\in H^{0}(C,\mathcal{L})$ comprising $f:C\to\mathbb{P}^{r}$ have the same order of vanishing for $j\in(S_{\circ}^{i})^{c}$. Therefore, twisting $\mathcal{L}$ down by this order of vanishing yields a point $y_{k^{\prime}}$ in the sequence $y=(y_{0},\ldots,y_{k})$ for which the corresponding coordinates are all non-zero and equal to each other, for the same reason as in Lemma 3.13. $\Diamond$
@@
The quantities $\deg(\mathcal{L}),\deg(D_{S})$ are taken to be those that are changing throughout the twisting algorithm. As for the last term, we somewhat abusively set the values of $(r-\#S^{i}_{\circ})$ to be equal to $r$ initially (as if all marked points $p_{i}$ are initially regular), so that the last summation is initially equal to rn-rn, and, by (1), $\Omega$ is initially equal to zero. Then, if $p_{i}$ is inactive, the summand $(r-\#S^{i}_{\circ})$ is removed after either step 4 or 5 (whichever is applied first). If $p_{i}$ is wild, then the value of $S^{i}_{\circ}$ is set to the correct one after step 6. If $p_{i}$ is regular, the summand $(r-\#S^{i}_{\circ})$ remains equal to $r$ throughout.
By Lemma 3.14, the final value of $\Omega$ is an upper bound for the expected relative dimension of $\tau^{\prime}$. Therefore, if $f_{\text{init}}\notin\mathbb{P}^{\circ}$, that is, at least one step of the algorithm is run, then the expected relative dimension of $\tau^{\prime}$ will be strictly negative, as needed.
@@ 3. Step 3
    First, decreasing $\deg(\mathcal{L})$ by $\alpha_{\text{tot}}-\alpha_{\max}$ decreases $\Omega$ by $(r+1)(\alpha_{\text{tot}}-\alpha_{\max})$, and replacing each $D_{\{j\}}$ by $D_{\{j\}}-\alpha_{j}p$ increases it by $(r-1)\alpha_{\text{tot}}$. Finally, the modification of $D_{[r+1]\backslash S^{m}}$ decreases $\Omega$ by $(\#S^{m}-1)(\alpha^{\prime}(m)-\alpha^{\prime}(m-1))$. We thus need to show that
    $2\alpha_{\text{tot}}+(r+1)\alpha_{\max}-\sum_{m=2}^{t}(\#S^{m}-1)(\alpha^{\prime}(m)-\alpha^{\prime}(m-1))<0.$
@@ 5. Step 5
  • if $p_{i}$ is wild, then step 6 decreases $\deg(\mathcal{L})$ and $\deg(D_{\{j\}})$ together by at least 1 for each $j\in S^{i}_{\circ}$, decreasing $\Omega$ by at least $2\cdot\#S^{i}_{\circ}$. Furthermore, the $(r-\#S^{i}_{\circ})$ is updated to the correct value (from r-r originally) in this step. In total, the value of $\Omega$ goes down at least by
    $2\cdot\#S^{i}_{\circ}+(r-\#S^{i}_{\circ})-r>0$
    in step 6. In step 7, $\Omega$ can again only decrease, by step 3. (Note that step $7$ does not affect the contribution of $(r-\#S^{i}_{\circ})$.)
@@
Therefore, the conclusion follows from Lemma 1.1. $\Diamond$
Proof of Proposition 3.2.
Lemmas 3.17 and 3.18 at long last imply that $V(x_{1},\ldots,x_{n})$ is contained in $\mathbb{P}^{\circ}$. Indeed, if this is not the case for some $f\in V(x_{1},\ldots,x_{n})$, then applying the twisting algorithm to $f$ yields a point on ${\mathcal{M}}_{g,n}(Y,\bar{\beta})$ moving in a family dominating the target of $\tau^{\prime}$, but this is a contradiction due to dimension reasons.
@@
It remains to check that any $f\in V(x_{1},\ldots,x_{n})$ has no non-trivial tangent vectors. It is elementary to check that such a tangent vector would give rise to a non-trivial relative tangent vector of the map $\tau:{\mathcal{M}}_{g,n}(X,\beta)\to{\mathcal{M}}_{g,n}\times X^{n}$ over a general point, which contradicts Lemma 1.1. $\Diamond$
@@ 3.2. Integral formula on $S$
In this section, we push forward the formula of Theorem 1.10 to obtain Theorem 1.11. Up to setting $k_{i}=0$ for $i>\ell$, we can assume that $\ell=r+1$.
@@
$\widetilde{\mathcal{E}}:=\mathcal{E}\otimes\mathcal{O}(N)=\bigoplus_{i=1}^{r+1}\nu_{*}(\mathcal{P}(-\mathcal{D}+N+\mathcal{D}_{i}))=\bigoplus_{i=1}^{r+1}\nu_{*}(\mathcal{P}(\Delta_{i})),$
@@
[…]
@@
where in the last equality we have pushed forward to $S$. It remains to compute c(~)c($\widetilde{\mathcal{E}}$); this will be an application of Grothendieck-Riemann-Roch.
@@
3.2.2. Computation of c(~)c($\widetilde{\mathcal{E}}$)
We now express c(~)c($\widetilde{\mathcal{E}}$) in terms of the classes defined above. Fix an index $i$ with $1\leq i\leq r+1$.
Proposition 3.20.
@@ 3.2.2. Computation of c(~)c($\widetilde{\mathcal{E}}$)
where in the second equality we used identity (i) in Lemma 3.19. Similarly, we have
@@
Here, in the second equality we have used identity (ii) in Lemma 3.19. After summing all the obtained contributions, we obtain the claim. $\Diamond$
Corollary 3.21.
@@
Substituting 3.21 into (21) completes the proof of Theorem 1.11, namely that
@@ 3.3. Specialization to genus 0
We now specialize Theorem 1.11 to $g=0$ to prove Theorem 1.12.
@@
Then, the class $\eta_{i}$ is the first Chern class of $\mathcal{O}_{\mathbb{P}^{k_{i}}}(1)$ and Theorem 1.11 gives:
@@
This concludes the proof of Theorem 1.12.
@@ 3.4. Specialization to $X={\mathsf{Bl}}_{q}(\mathbb{P}^{r})$
Theorem 1.11 also admits a reasonably elegant specialization to $X={\mathsf{Bl}}_{q}(\mathbb{P}^{r})$. We take now $k_{2}=\ldots=k_{r+1}=0$ and write $k=k_{1}$, $\eta=\eta_{1}$, and $\mathsf{E}=\mathsf{E}_{1}$, so that $\tau=\overline{\tau}_{2}=\cdots=\overline{\tau}_{r+1}$ and $x=\overline{x}_{2}=\cdots=\overline{x}_{r+1}$.
@@
$=\mathrm{Coeff}\left(A(\eta)(C(\eta)(1+\eta B(\eta))-\eta D(\eta)^{2})^{g};\eta^{k}\right)$
@@
From Lemma 3.22, we obtain the residue formula
@@
${\displaystyle{m+n\choose r}=\sum_{k=0}^{r}{m\choose k}{n\choose r-k}}\text{ for all }r,n,m\in\mathbb{Z}_{\geq 0}.$
Finally, replacing $a$ with gmg-m, we obtain Theorem 1.13.
@@ 4. Virtual counts for $X={\mathsf{Bl}}_{q}(\mathbb{P}^{r})$
In this section, we establish the virtual count of Theorem 1.14.
@@ 4.1. Preliminaries on $QH^{*}(X)$
Because $X$ is Fano, by [21, Proposition 2.2], we have
@@
as $\mathbb{Q}[\mathsf{C}]$-modules. Here, $\mathsf{C}\subseteq H_{2}(X,\mathbb{Z})$ is the cone of effective curves in $X$, and by [11, Proposition 4.1], we have
@@
are the only two primitive sets in the sense of [15, Definition 1.1] and, using [15, Theorem 1.2], from these two sets we respectively obtain the following two relations in $QH^{*}(X,\mathbb{Q})$.
@@
so the above collection spans $QH^{*}(X,\mathbb{Q})$. The fact that they are linearly independent follows from [15, Theorem 1.2]. $\Diamond$
@@
of $H^{*}(X,\mathbb{Q})$. We now express each of these classes in $QH^{*}(X,\mathbb{Q})$ as linear combinations of the basis elements of Lemma 4.1 using [15, Theorem 1.6]. In order for [15, Theorem 1.6] to apply, we need the following result.
Lemma 4.2.
@@
Immediate consequence of [15, Theorem 3.9]. $\Diamond$
Lemma 4.3.
@@
We start with $\mathsf{H}^{i}$. Let $\alpha_{i}$ be the cone generated by $v_{r},\ldots,v_{i+1}$ for $i=1,\ldots,r$. We have that $\alpha_{i}$ corresponds to a toric subvariety of class $\mathsf{H}^{r-i}$. Then, [15, Theorem 1.6] and the fact the only exceptional set special for $\alpha_{i}$ is the empty set yields the desired equality.